find angle which is 25degree less then 4 times its compliment?

Answer by checkley79(3341)

x+y=90
x=4y-25
4y-25+y=90
5y=90+25
5y=115
y=115/5
y=23 ans.
x+2=90
x=90-23
x=67 ans.